xx IDBinBH

IDxBinxBH

Erik Schnetter <schnetter@cct.lsu.edu>

November 24 2009

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11eosxprecision 10Scope: privatex01  REAL

11Description: precision of EOS calculations in EOSOmni

11 Range 10 x01 Default: 1.e-8
110:*  10any positive number

11

11eosxtable  10Scope: privatex01  STRING

11Description: Initial data EOS identifyer

11 Range 10 x01 Default: 2DxPolytrope
11              10Any valid EOS name in EOSOmni

11

11eosxtablexfilepath 10Scope: privatex01  STRING

11Description: Directory containing the EOS file used to generate the initial data

11 Range 10 x01 Default: (none)
11              10Directory name, default (empty string) is ’do not tell Lorene to overwrite info found in Lorene data’

11

11filename  10Scope: privatex01  STRING

11Description: Input file name containing LORENE data

11 Range 10 x01 Default: (none)
11              10Any filename, potentially including path

11

11recalculatexeps 10Scope: privatex01  BOOLEAN

11Description: Should eps be recalculated (default), or taken from Lorene data?

11              10  x01  Default: yes

11

11initialxhydro 10Scope: shared from HYDROBASEx01  KEYWORD

11Extends ranges: 11MeudonxBinxNS

10LORENE Binary Neutron Star initial data

11

7 Interfaces

General Implements: meudonxbinxns

Inherits: grid summationbyparts admbase hydrobase

Grid Variables

7.0.1 PUBLIC GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 initxeosxkey 10 initxeosxkey x01 compact110
11                 10  x01 description11Initial data eos handle for EOSxOmni
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11INT

11

8 Schedule This section lists all the variables which are assigned storage by thorn EinsteinInitialData/MeudonxBinxNS. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11x11Conditional:
11x 11initxeosxkey

11x 11x

11

Scheduled Functions

HydroBasexInitial (conditional) x meudonxbinxnsxinitialise x set up binary neutron star initial data x

11x11Language:11c
11 x 11Reads: 11grid::coordinates(everywhere)
11 x 11Type: 11function
11 x 11Writes: 11initxeosxkey(everywhere)
11 x 11x  11admbase::metric(everywhere)
11 x 11x  11admbase::shift(everywhere)
11 x 11x  11admbase::dtshift(everywhere)
11 x 11x  11admbase::curv(everywhere)
11 x 11x  11admbase::alp(everywhere)
11 x 11x  11admbase::dtalp(everywhere)
11 x 11x  11hydrobase::eps(everywhere)
11 x 11x  11hydrobase::rho(everywhere)
11 x 11x  11hydrobase::vel(everywhere)
11 x 11x  11hydrobase::press(everywhere)
11 x 11x  11hydrobase::yxe(everywhere)
11 x 11x  11hydrobase::temperature(everywhere)